OpenStudy (anonymous):

What is the equation of a line passing through (-8, -3) and having a slope of -4?

OpenStudy (hybrik):

You can do Y=-4X + B

OpenStudy (hybrik):

Plug in the points to get -3=-4(-8) +B

OpenStudy (hybrik):

Simplify that to find the y intercept

OpenStudy (anonymous):

would it be any one of these y = −4x + 20 y = 4x + 20 y = −4x – 35 y = 4x – 35 y = -4x + 4

OpenStudy (anonymous):

I think it's D

OpenStudy (hybrik):

Simplifing -3=-4(-8) + B will get you -35 = B, so you will get -4x-35=y

OpenStudy (anonymous):

ok so C then